REACTION DYNAMICS OF VIBRATIONALLY EXCITED CH_{3}D MOLECULES WITH CHLORINE

CHRISTOPHER J. ANNESLEY, \underline{ANDREW E. BERKE}, F. FLEMING CRIM, The University of Wisconsin - Madison Department of Chemistry, 1101 University Avenue, Madison, WI 53705.

State-resolved reactions of CH_{3}D molecules with Cl atoms probe the consumption and disposal of vibrational and translational energy. Using state-resolved detection it is possible to determine the relative effects of these on the reaction and gain a better understanding of the reactivity differences between modes.

In one recent experiment, we explored the effect of exciting either the combination of the C-H symmetric stretch and the C-D stretch (ν_{1} + ν_{2}), the combination of the C-D stretch and the C-H antisymmetric stretch (ν_{2} + ν_{4}), or the combination of the C-D stretch and the first overtone of the CH_{3} bend (ν_{2} + 2ν_{5}). The reaction dynamics largely follow a spectator picture in which the surviving bond retains its initial vibrational excitation. In at least 80% of the reactive encounters of vibrationally excited CH_{3}D with Cl, cleavage of the C-H bond produces CH_{2}D radicals with an excited C-D stretch, and cleavage of the C-D bond produces CH_{3} radicals with an excited C-H stretch. Deviations from the spectator picture seem to reflect mixing in the initially prepared eigenstates and, possibly, collisional coupling during the reaction.